OpenStudy (anonymous):

what is x^2 times radical(x+3)? and what is the domain?

jimthompson5910 (jim_thompson5910):

The expression that will restrict the domain is \[\large \sqrt{x+3}\] since you cannot take the square root of a negative number. So \[\large x+3\ge0\] which means that \[\large x\ge-3\]

OpenStudy (amistre64):

your radical needs to be indexed it youre going to write it like that

OpenStudy (anonymous):

is it possible to solve for x in this equation?

jimthompson5910 (jim_thompson5910):

you mean in the original expression or the inequality I just wrote out?

OpenStudy (anonymous):

original expression

jimthompson5910 (jim_thompson5910):

no you cannot solve for x in the original expression because there is no equal sign or inequality sign

OpenStudy (anonymous):

well the full question started with (h*g)(x)=___? and h=x^2 and g=radical x+3

OpenStudy (anonymous):

??

jimthompson5910 (jim_thompson5910):

is that h times g or is that h composes g?

OpenStudy (anonymous):

it is a circle in the question so i assume it is multiplication?

OpenStudy (anonymous):

okay its multiplication yes

jimthompson5910 (jim_thompson5910):

so it really looks like this \[\large (h\circ g)(x)\] ???

OpenStudy (anonymous):

yesss

jimthompson5910 (jim_thompson5910):

ok thx

jimthompson5910 (jim_thompson5910):

\[\large (h\circ g)(x)\] really means \[\large h(g(x))\], ie \[\large (h\circ g)(x)=h(g(x))\]

OpenStudy (anonymous):

ohh

jimthompson5910 (jim_thompson5910):

So start with h(x) and replace every x with g(x) to go from h(x)=x^2 to h(g(x))=(g(x))^2 Now replace g(x) in the right side to with sqrt(x+3) to get h(g(x))=(sqrt(x+3))^2 and now square the square root so it cancels out to give us h(g(x))=x+3 Note: this is assuming that x is nonnegative.

OpenStudy (anonymous):

-3 isnt the answer i tried on the website im doing this on

OpenStudy (anonymous):

oh wait the equation is the answer?

OpenStudy (anonymous):

got it!

jimthompson5910 (jim_thompson5910):

yes, \[\large (h\circ g)(x)=x+3\] where x is assumed to be nonnegative